OpenStudy (anonymous):

PLEASE HELP!~The tickets for a high school basketball game are $3.00 for students and $4.50 for adults. At Tuesday night's game, 108 tickets were sold for a total of $381.00. How many adult tickets were sold?

OpenStudy (noseboy908):

im on it

OpenStudy (noseboy908):

hold on

OpenStudy (noseboy908):

can you open up the image?

OpenStudy (anonymous):

No I couldn't:(

OpenStudy (anonymous):

I just get confused with all the numbers: I cant figure out the two equations:/

OpenStudy (noseboy908):

OK, to summarize what I wrote, the 2 equations are: "381=3s+4.50a" and "108=s+a" do you understand so far?

OpenStudy (anonymous):

yes

OpenStudy (noseboy908):

Multiply the 2nd eqn by 3 so that you get one variable that is the same in both eqns. to get "324=3s+3a" for the second eqn. understand so far?

OpenStudy (anonymous):

I was working on the substitution method but I understand this way!

OpenStudy (noseboy908):

so subtract the second eqn from the first one to get what?

OpenStudy (anonymous):

57 = 1.5a Then divide by 1.5 right?

OpenStudy (anonymous):

a = 38 So 38 adult tickets were sold, right?

OpenStudy (noseboy908):

no, don't confuse your variables. 38 student tickets were sold, not adult tickets. So, now take 108(total tickets sold) - 38. To get what? that number is your answer.

OpenStudy (anonymous):

ok so 108 - 38 = 70 adult tickets

OpenStudy (noseboy908):

correct my padawan. You are learning.

OpenStudy (anonymous):

lol Thanks:D

OpenStudy (noseboy908):

That is all I can teach you. Go forth and prosper.